Can I please have some help with this question. This is a problem given by my instructor. I have come up with a different answer every time. I am either leaving out a step or performing too many. THANKS!!!!
y/5 - 2/5 = 1/3 - y

Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

y/5 - 2/5 = 1/3 - y
-------------
Multiply thru by 15 to eliminate fractions
3y - 6 = 5 - 15y
Collect terms
18y - 6 = 5
18y = 11
y = 11/18

Answer by checkley77(12844)

y/5 - 2/5 = 1/3 - y
y/5+y=1/3+2/5
(Y+5Y)/5=(5+2*3)/15
6Y/5=(5+6)/15
6Y/5=11/15
6Y/5=11/15 CROSS MULTIPLY
15*6Y=11*5
90Y=55
Y=55/90
Y=11/18 ANSWER.
PROOF:
(11/18)/5-2/5=1/3-11/18
11/18*1/5-2/5=(6-11)/18
11/80-2/5=-5/18
(11-2*18)/80=-5/18
(11-36)/80=-5/18
-25/90=-5/18
-5/18=-5/18